Quillen Company is performing a post-audit of a project completed one year ago. The initial estimates were that the project would cost $227,000, would have a useful life of 9 years, zero salvage value, and would result in net annual cash flows of $43,700 per year. Now that the investment has been in operation for 1 year, revised figures indicate that it actually cost $231,000, will have a total useful life of 11 years (including the year just completed), and will produce net annual cash flows of $36,300 per year. Click here to view PV table.

Evaluate the success of the project. Assume a discount rate of 11%. (If the net present value is negative, use either a negative sign preceding the number eg -45 or parentheses eg (45). Round present value answers to 0 decimal places, e.g. 125. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Original estimate net present value $
Revised estimate net present value $


The project                                                           is is not a success.
Answer #1

As per the Case;

Let us evaluate each phase with the present values Annuity of the Cashflows.

So,

Initially the project started with 227,000

Present Values of Annuity @ 11% for 9 years is: 5.5370

Then, PV of Cash Flows=Net Annual Inflow X PV factor

46000X5.5370=254,702

Net Present Value=Present Value of Cashflows-Capital Investment

254,702-227,000=27,702

If the project is revised

Present Value of Annuity will be @11% for 11 is:6.2065

Then, PV of Cash Flows=Net Annual Inflow X PV factor

36,300X6.2065=225,296

Net Present Value=Present Value of Cashflows-Capital Investment

225,296-231,000=(5,704)

Hence, The result shows that initial investment is positive with the amount of 27,702 but the revised project has a negative investment of (5,704). Hence the project is NOT SUCCESSFUL.

Original Estimate Net Present Value: 254,702-227,000=27,702

Revised Estimate Net Present Value: 225,296-231000=(5,704)